# Sub-diffraction estimation, discrimination and learning of quantum states of light

### v1

@@ Abstract
The resolution of optical imaging is classically limited by the width of the point-spread function, which in turn is determined by the Rayleigh length. Recently, spatial-mode demultiplexing (SPADE) has been proposed as a method to achieve sub-Rayleigh estimation and discrimination of natural, incoherent sources. Here we show that SPADE is optimal in the broader context of machine learning. To this goal, we introduce a hybrid quantum-classical image classifier that achieves sub-Rayleigh resolution. The algorithm includes a quantum and a classical part. In the quantum part, a physical device (demultiplexer) is used to sort the transverse field, followed by mode-wise photon detection. This part of the algorithm implements a physical pre-processing of the quantum field that cannot be classically simulated without essentially reducing the signal-to-noise ratio. In the classical part of the algorithm, the collected data are fed into an artificial neural network for training and classification. As a case study, we classify images from the MNIST dataset after severe blurring due to diffraction. Our numerical experiments demonstrate the ability to learn highly blurred images that would be otherwise indistinguishable by direct imaging without the physical pre-processing of the quantum field.
@@ I Introduction
Improving the resolution of optical imaging is a long-standing goal with a broad impact on science and technology, from astronomical observations to biomedical applications. New schemes and methodologies are most commonly benchmarked against the well-known Rayleigh resolution criterion, which heuristically states that it is hard to resolve any detail smaller than the width of the point-spread function (PSF). The latter is of the order or the Rayleigh length $\mathrm{x_{R}}=\lambda D/R$, where $\lambda$ is the wavelength, $R$ the radius of the pupil of the optical system, and $D$ the distance to the object Rayleigh (1879). A number of approaches have been developed to circumvent the Rayleigh resolution limit. Among the most important examples we recall super-resolution fluorescence microscopy Hell and Wichmann (1994), which exploits non-linear optics to deactivate neighboring emitters, and the use of $N$-photon states, which have an effective wavelength that is $N$-times smaller Boto et al. (2000); Kok et al. (2001); Giovannetti et al. (2009).
@@
Contrary to common belief, the Rayleigh criterion does not represent by any means the fundamental limit to optical resolution. In fact, this criterion only applies to direct imaging (DI), i.e. pixel-by-pixel measurements of the field intensity. In general, information is also carried by the phase of the quantum electromagnetic field, and can be extracted by an optical quantum computer, suitably programmed for the task of performing imaging with maximum resolution. While a universal optical computer may require high-order nonlinear interaction, it has been proven that linear interferometry is optimal for such a task Tsang et al. (2016); Nair and Tsang (2016); Lupo and Pirandola (2016); Lupo et al. (2020). Relevant applications include the estimation and discrimination of natural, incoherent sources, in the regime of ultra-weak signals where the mean photon number per detected pulse is much below one and multiphoton events are suppressed.
To make our presentation more concrete, here we focus on a particular interferometric scheme, dubbed Hermite-Gaussain (HG) spatial-mode demultiplexing (SPADE) Tsang et al. (2016), where the optical field is sorted in its transverse components along the HG modes. A schematic representation is shown in Fig. 1. This can be achieved, e.g., by exploiting spatial light modulators Paúr et al. (2016, 2018); Zhou et al. (2019); Pushkina et al. (2021); Zhou et al. (2023); Frank et al. (2023), image inversion Tham et al. (2017), a photonic lantern Salit et al. (2020), or multi-plane light conversion Boucher et al. (2020); Santamaria et al. (2023); Rouvière et al. (2024); Santamaria et al. (2024). HG SPADE is well suited for optical systems characterised by a Gaussian PSF, or in the limit where the object is much smaller than the Rayleigh length. In the regime of ultra-weak signals, HG SPADE is the optimal measurement strategy that achieves the ultimate quantum limit for problems like parameter estimation Tsang et al. (2016); Lupo and Pirandola (2016); Lupo et al. (2020); Krovi (2021) and binary hypothesis testing Lu et al. (2018); Zhou and Jiang (2019); Huang and Lupo (2021); Grace and Guha (2022); Schlichtholz et al. (2024). Furthermore, SPADE may boost optical resolution even in the regime of bright sources Pushkina et al. (2021); Boucher et al. (2020); Santamaria et al. (2023); Frank et al. (2023) due to improved signal-to-noise ratio.
Figure 1: The left part of the figure shows a schematic representation of image formation through an optical imaging system in the far field, paraxial approximation, and in the Sub-Rayleigh regime. The right part of the figure schematically represents measurement of the field by HG SPADE. HG SPADE is a measurement strategy where the focused is processed by a spatial demultiplexer that sorts the transverse field in its Hermite-Gaussian component. Finally, each component is independently measured by photon counting. The figure shows only the lower modes $\text{HG}_{00}$, $\text{HG}_{01}$, $\text{HG}_{10}$. This is in contrast with direct imaging, where the focused image is directly measured with a camera by pixel-by-pixel photon detection.
I.1 Summary of results and relation with previous work
In this work we achieve two main goals. First, we shed light on why HG SPADE is an optimal measurement to solve estimation and discrimination problems for ultra-weak sources in the sub-Rayleigh regime. Second, we address the more general problem of learning quantum state of light, i.e., we consider a setting where minimal information is given about the source objects. We show the advantage of HG SPADE persists even in this more general setting.
The optimality of HG SPADE for parameter estimation has been shown by direct calculations and comparison with the ultimate quantum limit Tsang et al. (2016). The main theoretical tools are the Fisher information and quantum Fisher information, which bound the statistical error through the (quantum) Cramér-Rao bound Paris (2009); Sidhu and Kok (2020). Other works have shown that the optimal measurement can be constructed by taking derivative of the PSF Paúr et al. (2016), and that HG SPADE is optimal for estimating the moments of the intensity distribution of incoherent sources Zhou and Jiang (2019).
For hypothesis testing, the main theoretical tools are the Helstrom bound, the Chernoff bound, and the relative entropy. These quantities yield the optimal guessing probability and have been directly computed in recent works Grace and Guha (2022); Lu et al. (2018); Huang and Lupo (2021); Huang et al. (2023). In particular, Grace and Guha showed that the Chernoff bound can be expressed in terms of the second moments of the intensity distribution of the source objects, also showing by direct calculation the optimality of SPADE Grace and Guha (2022).
The above-mentioned works rigorously and comprehensively proved the optimality of HG SPADE and advantages with respect to DI, but somewhat lacked a simple and direct physical explanation for these results. In this work we fill this gap with a simple theory that gives a direct physical intuition of the optimality of HG SPADE. Our approach is based on expanding the state of a single photon in the sub-Rayleigh regime. This expansion shows the special role played by HG modes, from which optimality for hypothesis testing and estimation follows in a straightforward way.
Moving to the second objective of this work, we note that the problems of hypothesis testing and parameter estimation, which are those most commonly discussed in literature, refer to a somewhat contrived setting where abundant prior information is available about the scene that is observed, with one or at most a few unknown parameters. This is in contrast with most common use cases of imaging, where one has little or no prior information about the objects in the scene. For example, for the problem of parameter estimation, a commonly discussed setup is the one when one knows that there are two point-like emitters, and the goal is to estimate their spatial separation Tsang et al. (2016) (or variants of this problem Lupo et al. (2020); Krovi (2021)). In statistical hypothesis testing, two known sources are given with different geometry Grace and Guha (2022): for example a single emitter vs a pair of emitters, or two different QR codes. In summary, in all these examples the physical situation is well modeled and there is only one unknown parameter, which has continuous values in parameter estimation and discrete values in hypothesis testing (with extensions to a few parameters in multi-parameter estimation Řehaček et al. (2017); Yu and Prasad (2018); Napoli et al. (2019) and $M$-ary hypothesis test Grace and Guha (2022), respectively).
@@ I.1 Summary of results and relation with previous work
In this work we move a step forward in the direction of real-world applications of imaging, where one has only little prior information about the scene that is observed. Here we show that SPADE still outperforms DI in this more general setting. Our application can be viewed as a hybrid quantum-classical machine learning algorithm for image classification. The goal is to learn highly blurred images, generated by unknown objects sampled from a large dataset, using the data collected from a limited number of photon detection events. As a case study we use MNIST (Modified National Institute of Standards and Technology) database of handwritten digits. The physical part of our scheme is the application of HG SPADE, which implements a pre-processing of the quantum information encoded in the unknown state of a single photon. The classical part is a machine learning algorithm, which takes the measurement outputs of HG SPADE as input for training and classification
@@
Here we simulate the physical part of the scheme, including the source emission, light propagation, and detection by HG SPADE or DI, assuming imaging through a diffraction-limited optical system (e.g., a microscope or telescope) in the regime of ultra-weak signals. The simulated data are then fed into a machine learning algorithm. Our numerical experiments show that HG SPADE combined with machine learning allows us to classify images with accuracy much above what can be achieved by DI.
I.2 Structure of the paper
In Section II we introduce the physical model and the relevant assumptions of our work. In Section III we present a theory that elucidates the special role played by the HG modes, hence explaining why HG SPADE is the optimal measurement strategy for estimation and discrimination problems. In Section IV we illustrate our numerical experiments, the results obtained and the methodology employed, with the goal of assessing the accuracy achieved by HG SPADE in learning sub-Rayleigh objects in comparison to DI. Finally, in Section V we present our conclusions and outline potential future directions of research.
@@ II The physical model
Consider an extended source of incoherent, quasi-monochromatic light at wavelength $\lambda$. Such an object is characterised by its source intensity distribution. For example, later in our numerical experiments we choose objects from the MNIST dataset: grey-scale masks of $28\times 28$ pixels, with intensity $I_{xy}$, where $x,y$ is the pixel location in the object plane. Our physical model is schematically shown in Fig. 1.
@@
The object is observed using an optical imaging system, which in the far field and within the paraxial approximation is characterised by its PSF, denoted $T$ Goodman (2008). The PSF is the image of a point-like source. In particular, $T(x^{\prime}-Mx,y^{\prime}-My)$ is the amplitude of the field at position $x^{\prime},y^{\prime}$ in the image plane given that it was emitted at position $x,y$ in the object plane, where $M$ is the magnification factor (to simplify the notation, from now on we put $M=1$). The particular form of the PSF depends on the details of the optical system. As a concrete example, here we consider a Gaussian PSF,
@@
The width of the PSF quantifies the amount of diffraction in the optical system, and is of the order of the Rayleigh length, $\sigma\simeq\mathrm{x_{R}}$. The Gaussian PSF can be used to approximate the PSF arising from a circular pupil, especially if $\sigma$ is much larger than the size of the source Santamaria et al. (2023).
@@ II.1 Direct imaging
Equation (4) represents the state of the field in the image plane. This is the state that is measured to extract information about the object. DI consists of pixel-by-pixel photon detection in the image plane, which yields measurement outcomes sampled from the probability distribution
$p_{xy}^{\text{DI}}$ $=\varepsilon\sum_{x^{\prime}y^{\prime}}I_{x^{\prime}y^{\prime}}|\langle 1_{xy}|T_{x^{\prime}y^{\prime}}\rangle|^{2}$ (5)
$=\varepsilon\sum_{x^{\prime}y^{\prime}}I_{x^{\prime}y^{\prime}}|T(x-x^{\prime},y-y^{\prime})|^{2}\,,$ (6)
whereas no photon is detected with the complementary probability $1-\varepsilon$. This probability density is the convolution of the object intensity distribution with the squared PSF. For Gaussian PSF, this is a Gaussian convolution:
$p_{xy}^{\text{DI}}=\varepsilon\mathcal{N}\sum_{x^{\prime}y^{\prime}}I_{x^{\prime}y^{\prime}}\,e^{-\frac{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}{2\sigma^{2}}}\,.$ (7)
Finally, we consider a finite number $N$ of photons detected. Therefore, the quantity of interest for us is the relative frequency, denoted $f_{xy}^{\text{DI}}(N)$, which is obtained by sampling (with replacement) from the probability distribution in (7).
As an example, Fig. 2 shows the original MNIST object and the relative frequencies $f_{xy}^{\text{DI}}(N)$ obtained from the simulation of $N=10^{3}$ photon detection events, in correspondence of PSF width $\sigma=4$ and $\sigma=8$ (in units of number of pixels).
Figure 2: An example of MNIST object and the relative frequency from DI simulated for $N=10^{3}$ photon detection events. The PSF width $\sigma$ is given in number of pixels.
@@ II.2 Spatial-mode demultiplexing
Optics offers to the experimenter much more than DI. The optical field at the image plane can in fact be measured in other ways. Among all possible detection strategies, a special role is played by SPADE, which is implemented through a linear, passive interferometer. This implies that the single-photon state in Eq. (4) is processed without increasing the average photon number. In presence of loss, there is a non-zero probability that the photon is lost and the interferometer outputs the vacuum.
@@
and $H_{u}$ denotes the Hermite polynomials. Note that the width $\sigma$ matches that of the Gaussian PSF, which yields the optimal choice of measurement.
Let us denote as $b_{uv}^{\dagger}$ the operator that creates a photon in the output port $(u,v)$ of the interferometer, for $u,v=1,\dots,n$. Such an interferometer, when applied to single-photon states, is represented by the operator
@@
$p_{uv}^{\text{SPADE}}$ $=\varepsilon\sum_{xy}I_{xy}|\langle 1_{uv}|W|T_{xy}\rangle|^{2}$ (12)
$=\langle\Psi_{uv}|\rho|\Psi_{uv}\rangle\,,$ (13)
with $\rho$ as in Eq. (4). A device that demultiplexes the optical field in the lower-order HG modes is schematically represented in Fig. 1.
@@
III The special role played by the HG basis
To elucidate the special role played by the HG, we expand the state in the HG basis in the limit where the object size is much smaller than the Rayleigh length. This expansion will show that all relevant information is encoded in the lower HG modes.

### v2

# Machine learning with sub-diffraction resolution in the photon-counting regime

@@ Abstract
The resolution of optical imaging is classically limited by the width of the point-spread function, which in turn is determined by the Rayleigh length. Recently, spatial-mode demultiplexing (SPADE) has been proposed as a method to achieve sub-Rayleigh estimation and discrimination of natural, incoherent sources. Here we show that SPADE yields sub-diffraction resolution in the broader context of image classification. To achieve this goal, we outline a hybrid machine learning algorithm for image classification that includes a physical part and a computational part. The physical part implements a physical pre-processing of the optical field that cannot be simulated without essentially reducing the signal-to-noise ratio. In detail, a spatial-mode demultiplexer is used to sort the transverse field, followed by mode-wise photon detection. In the computational part, the collected data are fed into an artificial neural network for training and classification. As a case study, we classify images from the MNIST dataset after severe blurring due to diffraction. Our numerical experiments demonstrate the ability to classify highly blurred images that would be otherwise indistinguishable by direct imaging without the physical pre-processing of the optical field.
@@ I Introduction
Improving the resolution of optical imaging is a long-standing goal with a broad impact on science and technology, from astronomical observations to biomedical applications. New schemes and methodologies are most commonly benchmarked against the well-known Rayleigh resolution criterion, which heuristically states that it is hard to resolve any detail smaller than the width of the point-spread function (PSF). The latter is of the order or the Rayleigh length $\mathrm{x_{R}}=\lambda D/R$, where $\lambda$ is the wavelength, $R$ the radius of the pupil of the optical system, and $D$ the distance to the object Rayleigh1879 .
@@
Contrary to common belief, the Rayleigh criterion does not represent by any means the fundamental limit to optical resolution. In fact, this criterion only applies to direct imaging (DI), i.e. pixel-by-pixel measurements of the field intensity. In general, information is also carried by the phase of the quantum electromagnetic field, and can be extracted by an optical quantum computer, suitably programmed for the task of performing imaging with maximum resolution. While a universal optical computer may require high-order nonlinear interaction, it has been proven that linear interferometry is optimal for such a task TsangPRX , PhysRevLett.117.190801 , PhysRevLett.117.190802 , PhysRevLett.124.080503 . In particular, spatial-mode demultiplexing (SPADE) can achieve optimal performance in a variety of setting HKrovi , lu2018quantum , Modern , GracePRL , PhysRevLett.127.130502 , Schlichtholz:24 , PhysRevA.96.062107 , Prasad , astro22 .
The primary goal of this work is to address the problem of image classification, in the sub-Rayleigh regime and in the limit of ultra-weak incoherent sources, using data collected from a finite number of photon detection events. We focus on detection schemes that are derived from a particular interferometric scheme, dubbed Hermite-Gaussian (HG) SPADE TsangPRX , where the optical field is sorted in its transverse components along the HG modes paur2016achieving , Paur:18 , Zhou2019 , Lvovsky , Zhou:23 , Frank:23 , Tham2017 , Salit:20 , Boucher2020 , Santamaria22 , Treps23 , Amato23 . Here we consider both HG SPADE and its extensions where linear combinations of the HG modes are considered.
Until now, most applications of SPADE have been limited to the problems of parameter estimation and binary hypothesis testing. In the latter (see e.g. GracePRL , lu2018quantum , PhysRevLett.127.130502 , astro22 ) two quantum states, $\rho_{0}$ and $\rho_{1}$, are given, and the goal is to find an optimal measurement to discriminate them ($M$-ary discrimination was also considered GracePRL ). Here we address a different problem. We consider a large database of quantum states $\rho_{j,k}$, where $j=0,1$ labels two classes of states — we also study the multiclass scenario — and $k=1,\dots,K$ indicates the elements in each class. Our goal is to find an optimal measurement and post-processing algorithm to associate quantum states to the different classes. While work in this direction has been reported with bright signals and coherent detection Lvovsky , Frank:23 , to the best of our knowledge our contribution is the first attempt to address image classification in the photon-counting regime.
We remark that the two problems (hypothesis testing and classification) are related but different. It is easy to show that they are equivalent only in the single-shot scenario (where only one photon is detected) and otherwise inequivalent for $N>1$ photon detections 111 Consider the problem of $M$-ary classification of quantum states, $\rho_{j,k}$, where $j=1,\dots,M$ labels the classes, and $k=1,\dots,K$ labels the elements in each class. Consider a measurement modeled by the POVM elements $\Lambda_{m}$, for $m=1,\dots,M$. If only one copy of the unknown state is available, then the average probability of successful classification is $P_{s}=\frac{1}{MK}\sum_{k}\sum_{j}\mathrm{Tr}\left(\Lambda_{j}\rho_{j,k}\right)=\frac{1}{M}\sum_{j}\mathrm{Tr}\left(\Lambda_{j}\rho_{j}\right)\,,$ (1) where $\rho_{j}=\frac{1}{K}\sum_{k}\rho_{j,k}$ (2) denote the ensemble averages within each class. This shows that, in the single-copy scenario, the classification problem is equivalent to the problem of hypothesis testing defined on the average states $\rho_{j}$. In the multi-copy scenario, this equivalence is lost. In the multi-copy scenario we are allowed to measure $N$ copies of the unknown state using a POVM with elements $\Lambda_{m}^{(N)}$. The probability of successful classification is $P_{s}^{(N)}=\frac{1}{MK}\sum_{k}\sum_{j}\mathrm{Tr}\left(\Lambda_{j}^{(N)}\rho_{j,k}^{\otimes N}\right)$ (3) This is not the same as the probability of successful discrimination of the averages states since $\frac{1}{K}\sum_{k}\rho_{j,k}^{\otimes N}\neq\rho_{j}^{\otimes N}\,.$ (4) .
@@
As a case study we consider the MNIST (Modified National Institute of Standards and Technology) database of handwritten digits. Our application may be viewed as a hybrid machine learning algorithm for image classification. The physical part of our scheme is the application of SPADE, which implements a pre-processing of the information encoded in the unknown quantum state of a single photon. The computational part is a machine learning algorithm, which takes the measurement outcomes as inputs data for training and classification
@@
Here we simulate the physical part of the scheme, including the source emission, light propagation, and detection by SPADE, or DI for comparison, assuming imaging through a diffraction-limited optical system (e.g., a microscope or telescope) in the regime of ultra-weak signals. Samples of $N$ photon detection events are obtained through Monte Carlo simulation. The simulated data are then fed into a machine learning algorithm. Our numerical experiments show that HG SPADE (or its modifications) combined with machine learning allows us to classify images with accuracy much above what can be achieved by DI.
The manuscript develops as follows. In Section II we introduce the physical model and the relevant assumptions of our work. In Section III we present the expansion of a generic weak incoherent state in the HG basis. In Section IV we illustrate our numerical experiments, including results methodology, for both binary and multiclass classification. Finally, in Section V we present our conclusions and outline potential future directions of research.
@@ II The physical model
Consider an extended, nearly planar source of incoherent, quasi-monochromatic light at wavelength $\lambda$. Such an object is characterised by its source intensity distribution. For example, for our numerical experiments we choose objects from the MNIST dataset: grey-scale masks digits from 00 to $9$, of $28\times 28$ pixels, with intensity $I_{xy}$, where $x,y$ is the pixel location in the object plane.
@@
The object is observed using an optical imaging system, which in the far field and within the paraxial approximation is characterised by its PSF, denoted $T$ goodman2008introduction . The PSF is the image of a point-like source. In particular, $T(x^{\prime}-Mx,y^{\prime}-My)$ is the amplitude of the field at position $x^{\prime},y^{\prime}$ in the image plane given that it was emitted at position $x,y$ in the object plane, where $M$ is the magnification factor (to simplify the notation, from now on we put $M=1$). The particular form of the PSF depends on the details of the optical system. As a concrete example, here we consider a Gaussian PSF,
@@
The width of the PSF quantifies the effects of diffraction in the optical system, and is of the order of the Rayleigh length, $\sigma\simeq\mathrm{x_{R}}$. The Gaussian PSF can be used to approximate the PSF arising from a circular pupil, especially if $\sigma$ is much larger than the size of the source Santamaria22 .
@@
Figure 1: On the left: an example of MNIST object. On the right: the relative frequencies of photon detection from DI, obtained through Monte Carlo simulation with $N=10^{3}$ photon detection events, for $\sigma=4,8$. The PSF width $\sigma$ is given in number of pixels. The color bar indicates the range of empirical relative frequencies.
Figure 2: The left part of the figure shows a schematic representation of image formation through an optical imaging system in the far field, paraxial approximation, and in the sub-Rayleigh regime. The image is pixeleted to convey the fact that we consider only a finite number of photon detection events. The right part of the figure schematically represents a measurement of the field by HG SPADE. HG SPADE is a measurement strategy where the focused is processed by a spatial demultiplexer that sorts the transverse field in its Hermite-Gaussian component. Finally, each component is independently measured by photon counting. The figure shows only the lower modes $\text{HG}_{00}$, $\text{HG}_{01}$, $\text{HG}_{10}$. In our work we also consider linear combinations of the HG states as well as higher modes.
@@ II.1 Direct imaging
Equation (8) represents the state of the field in the image plane. This is the state that is measured to extract information about the object. DI consists of pixel-by-pixel photon detection in the image plane. Conditioned on detecting a photon, this yields measurement outcomes sampled from the probability distribution
$p_{x^{\prime}y^{\prime}}^{\text{DI}}$ $=\sum_{xy}I_{xy}|\langle 1_{x^{\prime}y^{\prime}}|T_{xy}\rangle|^{2}$ (9)
$=\sum_{xy}I_{xy}|T(x-x^{\prime},y-y^{\prime})|^{2}\,.$ (10)
This probability density is the convolution of the object intensity distribution with the squared PSF. For Gaussian PSF, this is a Gaussian convolution:
$p_{x^{\prime}y^{\prime}}^{\text{DI}}=\mathcal{N}\sum_{xy}I_{xy}\,e^{-\frac{(x^{\prime}-x)^{2}+(y^{\prime}-y)^{2}}{2\sigma^{2}}}\,.$ (11)
Finally, we consider a finite number $N$ of photons detected. Therefore, the quantity of interest for us is the relative frequency, denoted $f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)$, which is obtained by sampling (with replacement) from the probability distribution in (11).
As an example, Fig. 1 shows an image in the MNIST database and the relative frequencies $f_{x^{\prime}y^{\prime}}^{\text{DI}}(N)$ obtained from the Monte Carlo simulation of $N=10^{3}$ photon detection events, in correspondence of PSF width $\sigma=4$ and $\sigma=8$ (in units of number of pixels).
@@ II.2 Spatial-mode demultiplexing
Optics offers to the experimenter much more than DI. The optical field at the image plane can in fact be measured in other ways. Among all possible detection strategies, a special role is played by SPADE, which is implemented through a linear, passive interferometer — a schematic representation is shown in Fig. 2. This implies that the quantum state of light in Eq. (8) is processed without changing the photon number (though in the presence of loss there is a non-zero probability that the photon is lost).
@@
and $H_{u}$ denote the Hermite polynomials. Note that the width $\sigma$ matches that of the Gaussian PSF.
Let us denote as $b_{uv}^{\dagger}$ the operator that creates a photon in the output port $(u,v)$ of the interferometer, for $u,v=1,\dots,n$. Such an interferometer, when applied to single-photon states, is represented by the unitary operator
@@
$p_{uv}^{\text{SPADE}}$ $=\sum_{xy}I_{xy}|\langle 1_{uv}|W|T_{xy}\rangle|^{2}$ (16)
$=\frac{1}{\varepsilon}\langle\Psi_{uv}|\rho|\Psi_{uv}\rangle\,,$ (17)
with $\rho$ as in Eq. (8), and the factor $1/\varepsilon$ comes from conditioning on photon detection. A device that demultiplexes the optical field in the lower-order HG modes is schematically represented in Fig. 2.
@@
III Expansion in the HG basis
In our numerical experiment we will make use of the an expansion of a generic single-photon state in terms of the HG modes.